two types of candy sell for $2.50 per pound and $4.00 per pound. A store clerk is trying to make a 5 pound mixture worth 17.60. How much of each type of candy should be included in the mixture?
Answer by mananth(16946)

5 lb mixture = $17
1 lb = 17/5 = 3.4
Value ---------------- quantity
Candy I 2.5 ---------------- x pounds
Candy II 4 ---------------- 5 - x pounds
Mixture 3.40 ---------------- 5

2.5x+4(5-x)=3.4*5
2.5x+20-4x=17
2.5x-4x =17-20
-1.5x = -3
/-1.5
x=2 pounds Candy I
3 pounds Candy II